Discuss with colleagues and IITians

if 3 element p,q and crystallise in cubic lattice with p atom at corners and q atom at cube centre and r atom at centre of edges.what is the formula of compound?

aswathi gopal
14 Years agoGrade

2 Answers

Abhishek IITD
14 Years ago

contribution of atoms on the corners= (1/8) each. Hence p=(1/8)* 8=1

contribution of atoms at the centre- 1 each. Hence q= 1*1

contribution of atoms at the edge centre= (1/4) each. Hence r= (1/4)* 12= 3

 

thus the formula will ne pqr3.

plz approve the answer if u like it

pankaj kumar raman
14 Years ago

p=1/8 *8=1

q=1=1

r=1/4 *12=3

 

hence  it is  pqr3